A 100 kg mass drop 3 m, resulting in an increased volume in the cylinder of 0.002m^3. the weight and the piston maintain constant gage pressure of 100KPa. Determine the net work done by the gas on surroundings. Neglect all friction
1
Expert's answer
2020-11-09T06:51:09-0500

As per the given question,



Word done by the drops "W_1=-F\times d =-100\times 9.8\times 3J"

"W_1=-2940J"

Work done by the system "W_2=PA\times h =100000\times 0.002J =200J"

Hence net required work done "W=-2940+200 =-2740J"